OpenStudy (anonymous):

write 2 x-values at which the function f(x)=-2 cos 2x has a minimum. Please help me!!!

OpenStudy (anonymous):

max value of cosine is when \(x=0\) so min value of \(-2\cos(2x)\) is there as well one answer is \(x=0\)

OpenStudy (anonymous):

i am sorry i meant to say maximum values of cosine is 1 when \(x=0\)

OpenStudy (anonymous):

so one answer is \(x=0\) then \(2\pi\) also gives 1, so set \(2x=2\pi\) solve for \(x\) get \(x=\pi\)